OpenStudy (anonymous):

-x-y=1 find three points to solve the equation, then plot them on a graph

OpenStudy (anonymous):

if i am understanding this right this is really simple

OpenStudy (anonymous):

all you have to do is plug in x and y values that make the equation true

OpenStudy (anonymous):

okaii so can u give meh some help to break it down

OpenStudy (anonymous):

okay, what 2 numbers can you replace the x and the y with to equal 1

OpenStudy (anonymous):

1 and 2

OpenStudy (anonymous):

probably easier just to pick an \(x\) and find the \(y\) i pick \(x=3\) then \(-3-y=1\) so \(-y=4\) making \(y=-4\) now i have the point \((3,-4)\)

OpenStudy (anonymous):

or you could make it even easier by taking \[-x-y=1 \] and solve for \(y\) you get \[-y=x+1\] so \[y=-x-1\] now when you pick an \(x\) you already have \(y\) solved for i pick \(x=7\) so \(y=-7-1=-8\) giving me the point \((7,-8)\)